solve logic expression and write truth table for both

WILUVIIDouo 11 Р a. M=(AB)+(CD) b. P = (AC+BC)(A+C)
Answer #1
a.
M = AB+CD
A B C D M
F F F F F
F F F T F
F F T F F
F F T T T
F T F F F
F T F T F
F T T F F
F T T T T
T F F F F
T F F T F
T F T F F
T F T T T
T T F F T
T T F T T
T T T F T
T T T T T
b.
P = (AC+BC)(A+C)
P = ACA+ACC+BCA+BCC
P = AC+AC+BCA+BC
P = AC+BCA+BC
P = (AC)(1+B)+BC
P = (AC)(1)+BC
P = AC+BC
P = (A+B)C
A B C P
F F F F
F F T F
F T F F
F T T T
T F F F
T F T T
T T F F
T T T T